OpenStudy (anonymous):

a magician decides to use an electric field to float a light plastic ball in the air . the ball has a mass of 15 grams and can be given a charge +0.5 coloumbs. a) calculate the weight of the ball b) determine the strength and direction of external electric field that the magician would require to float the ball plz plz help

OpenStudy (ujjwal):

Can you calculate the weight of the ball?

OpenStudy (kainui):

So F=ma and F=qE right? In order for an object to be still, forces must be in equilibrium, right?

OpenStudy (anonymous):

Well, since the charge on the ball is positive, the direction of electrostatic force will coincide with the direction of the electric field. So if the force of gravity cancels with the electrostatic force, the forces must be equal in magnitude but in opposite directions. Can you figure that out?

OpenStudy (ghazi):

both the forces should be balanced and in opposite direction therefore m*g= q*E............therefore 0.0015kg* 10= 0.5*E...now solve for E